OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = quantity x minus eight divided by quantity x plus seven. and g(x) = quantity negative seven x minus eight divided by quantity x minus one.

OpenStudy (anonymous):

Consider checking the other copy of this question.

OpenStudy (anonymous):

The other question was typed wrong

OpenStudy (anonymous):

Okay, well is this what you're saying? \[\large \begin{cases}f(x)=\dfrac{x-8}{x+7}\\\\\\ g(x)=\dfrac{-7x-8}{x-1}\end{cases}\]